OpenStudy (anonymous):

Chang has 2 shirts: a white one and a black one. He also has 2 pairs of pants, one blue and one tan. What is the probability, if Chang gets dressed in the dark, that he winds up wearing the white shirt and tan pants? Show your work.

OpenStudy (anonymous):

Wouldnt it be white and tan / black and blue = 2/4=1/2

OpenStudy (anonymous):

@jim_thompson5910 sorry I keep calling you for this I feel bad lol

OpenStudy (anonymous):

@kidrah69

OpenStudy (anonymous):

hi i can help you

OpenStudy (anonymous):

Thanks am I correct?

OpenStudy (anonymous):

Yes you are correct :)

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

call me for any question you have because i can answer it.